OpenStudy (anonymous):

Bozo the clown brings a total of 20 balloons to Jake ’ s birthday party. There are 7 yellow, 2 red , and 1 1 green balloons . He reaches into his pocket and randomly selects a color to make a balloon animal. If he pulls a yellow balloon first, what is the probability of the next color to be selected randomly will also be yellow?

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

First, add up all the numbers. What do you get?

OpenStudy (anonymous):

no you dont add all the numbers togehter

OpenStudy (anonymous):

Yeah and then put how much yellow over i9t for a fraction.

OpenStudy (anonymous):

but they took one yellow from there

OpenStudy (anonymous):

well the answer is how many yellow balloons minus one yellow balloon equals blank over total amount then see if there is a common number to be divided by

OpenStudy (anonymous):

so its 6 yellow

OpenStudy (anonymous):

by the total amount i mean how many balloons he has all together. yes he has 6 yellow and it is over 20 so divide both by 2 and you get ur answer

OpenStudy (michele_laino):

Hint: the yellow balloons, after the first selection, are 6, so the total balloon are 20-1=19. Now the requested probability is equal to the ratio: favorable cases/ possible cases

OpenStudy (anonymous):

long problem short u get 3/10

OpenStudy (anonymous):

so it would be 6/19 @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ok so i divide that

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I think that it is not necessary, namely p=6/19 is a good answer

OpenStudy (anonymous):

wait what that cant be the answer 6/19

OpenStudy (anonymous):

yeah it can. Probability can be a fraction.

OpenStudy (michele_laino):

furthermore, note that, the total probability, namely the probability to get two yellow balloons, at the first two selections, is equal to the product of the single probabilities, namely: total probability = (7/20)*(6/19)=...?

OpenStudy (anonymous):

so what is the answer?

OpenStudy (michele_laino):

the answer is: the probability of the second selection, is 6/19 whereas the total probability to get two yellow balloons consecutively, namely after two selections, is given by the subsequent product: (7/20)*(6/19)

OpenStudy (michele_laino):

in general the probability is expressed as a ratio, or fraction

OpenStudy (michele_laino):

so, the first probability, is 6/19, whereas the second probability or total probability, is: (7/20)*(6/19)=42/380=21/190